Solve the following equation:
square root of x-1=3

Answer by wuwei96815(245)

If I understand the problem correctly the easiest way to solve it is to square both sides of the equation.

Then you are left with: x-1 = 9. Solving for x we get x = 9+1 = 10.

I hope that is right.